Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: C configuration OST-driven science observing; ongoing testing for 3-bit, switched power, etc. Brent Carlson visiting (focused on WIDAR connectivity and robustness). Planned power outage at the end of the week (starting 0700 Friday through Saturday as needed). Wiki madness - staff wiki to public wiki migration - will send out updated link.
  • Key dates:
  • Staff Travel:
    • Perley: 23 Sep - 19 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • EA14 final integration (collimation checkout last night); EA24 headed in for electronics overhaul - out by next week.
      • Note problems with EA27 (50-100 ns jumps) for several months; being investigated.
      • Antennas with 3-bit samplers: EA12, EA15, EA22, EA28
    • Links:
  • WIDAR (Rupen):
    • Phased array testing (Adam/Joe) very successful (See: https://bugs.nrao.edu/browse/ECSV-100). Brent visiting. Plans for this week: switched power testing; CRM time; 3-bit follow-up; EA15's sampler brought back to the labs.
      • Time averaging test data taken last night (after solving OPT issue).
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • Notes: MCAF update made last Wednesday (no known issues); OPT 1.06 on webtest; meeting today to discuss testing. PST handling preparation for next cycle.
  • CASA Update (new functionality available/issues/enhancements):Release delayed for some performance enhancement issues.
  • AIPS Update (new functionality available/issues/enhancements): Ongoing work with switched power. Able to edit tables and plot Tsys or Tcal. Added to SNEDT - can flag based on uvdata (to get rid of off-source data). TSMO provides clipping and smoothing. SPLIT now handles sources with multiple calcodes. FGPLOT can handle any # of baselines. FRING does RL delay, can ask to fit 1st half of IFs together (AC vs BD fit); seems to help.
  • SDM Update: Updated document. Links.

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

  • CBE update testing
  • Switched power testing
  • CBE averaging sign-off tests (also pending response from first RSRO projects).
  • Next OSRO cycle testing (delayed from last week)
  • Science and...
  • Interrupt status for additional 3-bit testing

Commissioning Reports

Planetary: Butler: Pluto observations

  • Easily detected (400 micro Jy) but not resolved with Charon; large amount of flagged data (later understood due to a non-optimal LST start time range and low elevation).

Calibration: Dhawan: Switched power testing

  • Using OBIT bdf conversion to AIPS format. Still exploring but using new tools to flag on Tsys data (more next week).

Demo science: Fomalont: Fornax calibration and imaging results

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)

Commissioning Test Planning

  • OPT: webtest of 1.06 capabilities/bug fixes (*); ongoing (Lorant, Michael, Joe)
  • Calibration: Switched power on/off 1dB attenuators. Data taken: 10/18
  • Array: RF switch tests with rapid monitoring (as possible but likely not this week).
  • Correlator: CBE averaging with restricted nodes
  • Correlator: Quadrant 2 tests; CM update test.
  • Calibration: Switched power tracking over several airmasses.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed until Rick's return
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Spectral Line: Momjian: Zeeman test follow-upHigh frequency/low frequency results.

RSRO/ECSO Reports

10B-104: Kellerman/Kimble: First results: fair amount of interference seen. Surprised to see bandpass variations over a single subband (stability was noted as the key); more later.

AL746: Leroy: First results: problems on a subset of antennas; still being looked at.

10B-209: Hallinan: First results - not this week.

Other

  • Satoko: Also saw lots of flagged data; will send report to be rolled in.
  • Urvashi seeing missing subbands; will report.

Other updates:

  • AC982: Seeing delay jitter on EA08 on all IFs, subbands, etc (from August); note for C-band observations, the inclusion of the interim Rxs complicated the reduction (in a non-closing way); flagging these antennas is needed.

  • Notes:
    • Current Residents
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: Juergen Ott); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs
      • Jill Knapp (POC: Rupen/Claussen); area(s): recirculation; IRC+10216 demo science; 10B-200;10B-186. Welcome again.

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Commissioning Group updates

AOB

  • Brent talk on proposal for SKA correlator; next Thursday for discussion on Zeeman observations (Momjian).
  • Note: Solar observing discussions this week with Tim Bastian and with CCB.

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 2010-10-19
Topic attachments
I Attachment Action Size Date Who Comment
fornax_report-1.pdfpdf fornax_report-1.pdf manage 1 MB 2010-10-26 - 11:33 JosephMcMullin  
st_keyhole_neb_med_cropped.jpgjpg st_keyhole_neb_med_cropped.jpg manage 137 K 2010-10-19 - 11:56 PatrickMurphy Test, please ignore.
Topic revision: r4 - 2010-10-27, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback